# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ 4 Application to smooth and strongly dissipative potentials
Let $V$ be a potential that is $L$-smooth, $\alpha$-strongly convex outside some ball of radius $R$. Then, denoting $\mu\propto e^{-V}$ and $\kappa=L/\alpha$, it holds for any critical point $x^{*}$ that
$\text{Var}_{\mu}(X)\leq\mathbb{E}_{\mu}[\|X-x^{*}\|^{2}]\leq 8R^{2}\left(\kappa+1\right)^{2}+\frac{4d}{\alpha}\,.$
The proof is deferred to Appendix B.1.
Remark 15
We note that Ma et al., [31] and Chehab et al., [9] also derived bounds for the second moment of smooth and (weakly) dissipative potentials which a fortiori provides bounds for the variance. Yet the former suffers an exponential dependence in $LR^{2}$ and the latter implicitly depends on the the gradient of the potential at 00 which would have induced a dependence in $x$ when applied to $q_{t,x}$.
This upper-bound combined with Proposition 9 and Proposition 13 yields that the forward process $p_{t}^{V}$ is uniformly smooth along the trajectory.
Corollary 16
Assume $V$ is $C^{2}$, $L$-smooth and $\alpha$-strongly convex outside some ball of radius $R$ Then, denoting the condition number $\kappa=L/\alpha$, it holds that
$\|\nabla^{2}\log(p_{t}^{V})\|\leq 4(L+1)^{2}\left(8R^{2}\alpha(\kappa+1)^{2}+\frac{4d}{\alpha}+1\right)\,.$
The proof can be found in Appendix B.2. Note that numerous works in the reverse diffusion literature make the a priori assumption that the forward process is $L$-log-smooth [12, 24, 25]. The result above provides a non-trivial setting where this assumption does hold.
Remark 17
Note that our bounds extend the results of Mikulincer and Shenfeld, [36] and could be used in a similar fashion to recover estimates on the Log-Sobolev constants of strongly dissipative potentials. Yet, because of our residual dependence in $d$, we expect our resulting bounds to worsen those found in Ma et al., [31].
In order to apply our Theorem 11, it remains now to bound the second moment $m_{2}(\mu)$, the gradients of the forward process $\|\nabla\log(p_{t}^{V})\|,\|\nabla\log(p_{t}^{2V})\|$ and the density $\mu$ at 00 in the case where $V$ is $L$-smooth and $\alpha$-strongly convex outside some ball $B_{R}(m)$. In what follows, to ease notations, we shall assume that $\nabla V(0)=0$. Note that this extra assumption is benign since we can always compute any critical point $x^{*}$ up to arbitrary precision in polynomial time [23] and then sample from $\tilde{V}=V(\cdot+x^{*})$. Using Proposition 14, this extra assumption readily implies $m_{2}(\mu)\leq 8R^{2}\left(\kappa+1\right)^{2}+\frac{4d}{\alpha}$. We provide in Appendix B.3 polynomial bounds on the remaining quantities under Assumptions 2-3 which allows us to prove that strongly dissipative distributions can be sampled from in polynomial time in fixed dimension.
Corollary 18
Under Assumptions 2-3 and if we further assume $\nabla V(0)=0$, there exists a stochastic algorithm to sample from $\mu\propto e^{-V}$ that outputs a sample $Y\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ that runs in (deterministic) time $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$ where $\lesssim$ hides a universal constant and log factors in $d,L,\alpha,R$.
The proof is left in Appendix B.4. We summarize in Table 1 how our bounds compare to those of previous works after plugging our estimate of $m_{2}(\mu)$ in the smooth, strongly dissipative case. Assuming that our estimate is tight, we observe that while our bound exhibits a worse dependence in $\epsilon^{-1},d$ than in Ma et al., [31] and in He et al., [22], we provide an exponential improvement with respect to $R,L$ over previous works where we lowered the complexity to polynomial. We provide in Appendix D a 2D simulation that compares standard ULA to our algorithm.
Algorithm Complexity
ULA [31] $O(e^{32LR^{2}}\kappa^{2}\frac{d}{\epsilon^{2}})$
RD + ULA [25] $e^{O(L^{3}\log^{3}(L^{2}(R^{2}\kappa^{2}+d/\alpha)/\epsilon))}$
RD + Rejection Sampling [22] $O(L^{d/2}\epsilon^{-d}e^{L\|x^{*}\|^{2}+L^{2}(\kappa^{2}R^{2}+d/\alpha)})$
RD + Self-normalized IS (ours) $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$
Table 1: Complexity of sampling algorithms in the smooth, strongly dissipative case. We denoted $x^{*}$ the global minimum of $V$.
5 Conclusion
In this article, we pushed one step further the reduction from sampling to intermediate score estimation initiated over the past two years by Chen et al., 2023b [12], Chen et al., 2023a [10], Conforti et al., [16], Benton et al., [5]. Using our self-normalized estimator of the scores, we proved that in fixed dimension, the sampling problem could be further reduced to a regularity problem along the forward OU process. We then applied this result to derive polynomial time guarantees for the task of sampling from a log-smooth, strongly dissipative distribution thus providing an exponential improvement over previously known bounds. Yet we believe that our contribution could pave the way for numerous other non-log concave sampling problems in low dimensions. In future works, we plan to apply our results to Gaussian mixtures which are not fully covered by the strongly dissipative framework.
Acknowledgements
This work was funded by ANR JCJC WOS. We thank Pierre Monmarché for his help in proving Proposition 14 and Lemma 27.
Appendix A Proofs of Section 3
A.1 Identities of the forward process
Before starting our proofs, we recall useful identities on the density and the log hessian of the forward process known as the Tweedie’s formulas [39].
@@ A.1 Identities of the forward process
$p_{t}^{V}(z)=\frac{e^{td}}{Z^{V}}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]\,,$ (6)
@@
$\nabla^{2}\log(p_{t}^{V})(z)=\frac{\mathbb{E}[Y_{t}Y_{t}^{\top}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})^{2}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}-\frac{I_{d}}{(1-e^{-2t})}-(\nabla\log(p_{t}^{V})(z)))(\nabla\log(p_{t}^{V})(z))^{\top}\,.$ (7)
@@
[…]
@@
[…]
@@
A.2 Proof of Proposition 6
@@
$\mathbb{E}[\|\hat{N}(z)-N(z)\|^{2}]=\frac{1}{n}\mathbb{E}\left[\left\|\frac{Y_{t}e^{-V((z-Y_{t})e^{t})}}{1-e^{-2t}}-N(z)\right\|^{2}\right]\leq\frac{1}{n}\mathbb{E}\left[\frac{\|Y_{t}\|^{2}e^{-2V((z-Y_{t})e^{t})}}{(1-e^{-2t})^{2}}\right]\,.$
@@
$\mathbb{E}[(\hat{D}(z)-D(z))^{2}]=\frac{1}{n}\left(\mathbb{E}[e^{-2V((z-Y_{t})e^{t})}]-\mathbb{E}[e^{-2((z-Y_{t})e^{t})}]^{2}\right)\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$ (8)
A.3 Proof of Proposition 7
@@
To ease notations, we shall drop the dependence in $z$ throughout the proof. Define the event $A=(\hat{D}\geq\eta D)\cap(\|\hat{N}\|\leq\kappa\|N\|)$ where $\eta\leq 1,\kappa\geq 1$ are positive scalars to be chosen later. We start to decompose the quadratic error as:
@@
$\begin{cases}&\nabla\theta(x,p)=\frac{-2}{p}\left(\frac{N}{D}-\frac{x}{p},\left\|\frac{x}{p}\right\|^{2}-\left\langle\frac{x}{p},\frac{N}{D}\right\rangle\right)\,,\\ &\nabla^{2}\theta(x,p)=\frac{-2}{p^{2}}\begin{pmatrix}-I_{d}&(\frac{2x}{p}-\frac{N}{D})^{\top}\\ (\frac{2x}{p}-\frac{N}{D})&-3\left\|\frac{x}{p}\right\|^{2}+2\langle\frac{x}{p},\frac{N}{D}\rangle\end{pmatrix}\,.\end{cases}$
We thus make a Taylor expansion of order $2$ of $\theta(\hat{N},\hat{D})$ around $(N,D)$: there exists (a random) $\hat{t}\in[0,1]$ such that
@@
$\theta(\hat{N},\hat{D})$ $=\frac{1}{\hat{D}_{\hat{t}}^{2}}\bigg{(}\|\hat{N}-N\|^{2}-2(\langle 2\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}}-\frac{N}{D},\hat{N}-N\rangle(\hat{D}-D))+\left(3\frac{\|\hat{N}_{\hat{t}}\|^{2}}{\hat{D}_{\hat{t}}^{2}}-2\langle\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}},\frac{N}{D}\rangle\right)(\hat{D}-D)^{2}\bigg{)}$
@@
Hence, almost surely over $A$
@@
Let us now handle the quadratic error of the estimator on the complementary $\bar{A}$. We have, using Young’s inequality $||a-b||^{2}\leq 2(||a||^{2}+||b||^{2})$,
@@
Now, recall that $X_{i}=||y_{i}||^{2}(1-e^{-2t})^{-1}$ are $n$ independent variables such that for all $i$, $X_{i}\sim\chi^{2}(d)$. Using Hölder inequality for some $p\geq 1$, the second term can be upper-bounded as
@@
We now upper bound the probability of the event $\bar{A}=(\hat{D}<\eta D)\cup(\|\hat{N}\|>\kappa\|N\|)$. By Tchebychev inequality, using $\eta<1$, it holds that
@@
Similarly, recalling that $||N||=D||\nabla\log(p_{t}^{V})||$, we have
$\mathbb{P}(\|\hat{N}\|>\kappa\|N\|)$ $\leq\frac{\mathbb{E}[||\hat{N}-N||^{2}]}{||N||^{2}(\kappa-1)^{2}}$
@@
Combining this with the bound Eq. A.3 eventually yields

### v3

@@ B.4 Discussion on  Lytras and Mertikopoulos [2025]
Let $\mu=1/2\mathcal{N}(c,\lambda_{max}I_{d})+1/2\mathcal{N}(-c,\lambda_{min}I_{d})$ with $\lambda_{max}\geq\lambda_{min}>0$ and $\|c\|>0$. It holds that
$C_{P}(\mu)\geq\frac{\|c\|^{2}e^{\frac{\|c\|^{2}}{2\lambda_{max}}}}{2}\,.$
Proof.
Denoting $u=c/\|c\|$ and defining $g:\mathbb{R}\to\mathbb{R}$ as
[…]
consider the test function $f(x)=g(x^{\top}u)$. Denoting $\mu_{1}=\mathcal{N}(c,\lambda_{max}I_{d})$, $\mu_{2}=\mathcal{N}(-c,\lambda_{min}I_{d})$, we start to upper-bound the Dirichlet energy $I=\int\|\nabla f\|^{2}\mathrm{d}\mu$:
$I=\frac{2}{\|c\|^{2}}\big(\mathbb{P}_{X\sim\mu_{1}}\big[-\frac{\|c\|}{2}\leq X^{\top}u\leq\frac{\|c\|}{2}\big]+\mathbb{P}_{X\sim\mu_{2}}\big[-\frac{\|c\|}{2}\leq X^{\top}u\leq\frac{\|c\|}{2}\big]\big)\,.$
Now remark that for $X\sim\mu_{i}$, the random variable $X^{\top}u$ is also a one dimensional gaussian: if $X\sim\mu_{1}$, it holds that $X^{\top}u\sim\mathcal{N}(\|c\|,\lambda_{max})$ and if $X\sim\mu_{2}$, it holds that $X^{\top}u\sim\mathcal{N}(-\|c\|,\lambda_{min})$. In particular, the Dirichlet energy re-writes as
$I=\frac{2}{\|c\|^{2}}\big(\Phi(\frac{-\|c\|}{2\sqrt{\lambda_{max}}})-\Phi(\frac{-3\|c\|}{2\sqrt{\lambda_{max}}})+\Phi(\frac{-\|c\|}{2\sqrt{\lambda_{min}}})-\Phi(\frac{-3\|c\|}{2\sqrt{\lambda_{min}}})\big)\,$,
where $\Phi$ is the cumulative density function of the standard Gaussian. In particular, $I$ is upper-bounded as
$I\leq\frac{4\Phi(\frac{-\|c\|}{2\sqrt{\lambda_{max}}})}{\|c\|^{2}}=\frac{2}{\|c\|^{2}}(1-\text{erf}(\frac{\|c\|}{\sqrt{2\lambda_{max}}}))\leq\frac{2e^{-\frac{\|c\|^{2}}{2\lambda_{max}}}}{\|c\|^{2}}\,.$
Now there remains to lower-bound the variance term. The variance reads
$\text{Var}_{\mu}(f)=\int f^{2}\mathrm{d}\mu=1/2(\mathbb{E}_{\mu_{1}}[f^{2}(X)]+\mathbb{E}_{\mu_{2}}[f^{2}(X)])\,.$
The first term can be lower-bounded as
$\mathbb{E}_{\mu_{1}}[f^{2}(X)]$ $\geq\mathbb{P}_{X\sim\mu_{1}}[X^{\top}\mu\geq\frac{\|c\|}{2}]\,$,
$=\Phi(\frac{\|c\|}{2\sqrt{\lambda_{\max}}})\,$,
$\geq 1/2\,.$
Conversely, it holds that $\mathbb{E}_{\mu_{2}}[f^{2}(X)]\geq 1/2$ and a fortiori $\text{Var}_{\mu}(f)\geq 1$. Hence we recover that
$\frac{\text{Var}_{\mu}(f)}{I}\geq\frac{\|c\|^{2}e^{\frac{\|c\|^{2}}{2\lambda_{max}}}}{2}\,$,
which concludes the proof. ∎
B.5 Discussion on Agapiou et al. [2017]
In Agapiou et al. [2017], for $\nu$ dominated by $\pi$, denoting $g$ the unnormalised density ratio
$\frac{\mathrm{d}\nu}{\mathrm{d}\pi}(u)\coloneqq\frac{g(u)}{\int g(u)\mathrm{d}\pi(u)}\,$,
the authors derived error bounds for self-normalised estimators: for any test function $\phi:\mathbb{R}^{d}\to\mathbb{R}$, the expectation of $\phi$ under $\nu$, denoted $\nu(\phi)$, is estimated via
$\nu^{n}(\phi)=\frac{\sum_{i=1}^{n}\phi(x_{i})g(x_{i})}{\sum_{i=1}^{n}g(x_{i})}\,$,
with $(x_{i})$ $n$-iid samples drawn from $\pi$. In Theorem 2.3, they provide the following quadratic error bound:
$\mathbb{E}[(\nu(\phi)-\nu^{n}(\phi))^{2}]\lesssim\frac{3\pi(|\phi g|^{2d})^{1/d}m_{2e}(g)^{1/e}}{n\pi(g)^{4}}\,$, (22)
with $m_{t}(h)=\pi(|h(\cdot)-\pi(h)|^{t})$ and $d,e\in(1,+\infty[$ such that $1/d+1/e=1$. Now recall that, denoting $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$ the score along the forward is given by
$\nabla\log(p_{t}^{V})(z)$ $=\frac{-1}{1-e^{-2t}}\frac{\mathbb{E}[Y_{t}e^{-V(e^{t}(z-Y_{t}))}]}{\mathbb{E}[e^{-V(e^{t}(z-Y_{t}))}]}\,$,
$=\frac{-1}{1-e^{-2t}}\frac{\int ye^{-V(e^{t}(z-y))}e^{-\frac{\|y\|^{2}}{2(1-e^{-2t})}}\mathrm{d}y}{\int e^{-V(e^{t}(z-y))}e^{-\frac{\|y\|^{2}}{2(1-e^{-2t})}}\mathrm{d}y}\,$,
$=\frac{-1}{1-e^{-2t}}\mathbb{E}_{Z\sim\nu}[Z]\,$,
where $\nu(x)\propto e^{-V(e^{t}(z-x))}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}$. In particular, denoting $\pi(x)\propto e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}$, the quadratic error of the self-normalized estimator reads
$\|\hat{s}_{t,n}(z)-\nabla\log p_{t}^{V}(z)\|^{2}$ $=\frac{1}{(1-e^{-2t})^{2}}\biggl\|\frac{\sum_{i=1}^{n}y_{i}g(y_{i})}{\sum_{i=1}^{n}g(y_{i})}-\mathbb{E}_{Z\sim\nu}[Z]\biggr\|^{2}\,$,
$=\frac{1}{(1-e^{-2t})^{2}}\sum_{k=1}^{d}\biggl(\frac{\sum_{i=1}^{n}\phi_{k}(y_{i})g(y_{i})}{\sum_{i=1}^{n}g(y_{i})}-\mathbb{E}_{Z\sim\nu}[\phi_{k}(Z)]\biggr)^{2}\,.$
with $g=\mathrm{d}\nu/\mathrm{d}\pi$, $(y_{i})$ $n$-iid samples drawn from $\pi$ and $\phi_{k}:\mathbb{R}^{d}\to\mathbb{R}$ the projector on the $k$-th coordinate. Hence, the bound Eq. 22 of Agapiou et al. [2017] provides the following upper-bound
$\|\hat{s}_{t,n}(z)-\nabla\log p_{t}^{V}(z)\|^{2}\leq\sum_{k=1}^{d}\frac{3\pi(|\phi_{k}g|^{2d})^{1/d}m_{2e}(g)^{1/e}}{n\pi(g)^{4}}\,.$
We show in the next paragraph that, when integrated against $p_{t}^{V}$, this upper-bound is vacuous. By Jensen’s inequality, it holds that
$\begin{cases}&m_{2e}(g)^{1/(2e)}\geq m_{2}(g)^{1/2}\,,\\ &\pi(|\phi_{k}g|^{2d})^{1/(2d)}\geq\pi(|\phi_{k}g|^{2})^{1/2}\,.\end{cases}$
Hence, the previous right-hand side of is lower-bounded as
$\sum_{k=1}^{d}\frac{3\pi(|\phi_{k}g|^{2d})^{1/d}m_{2e}(g)^{1/e}}{n\pi(g)^{4}}\geq\sum_{k=1}^{d}\frac{3\pi(|\phi_{k}g|^{2})m_{2}(g)}{n\pi(g)^{4}}=\frac{3m_{2}(g)}{n\pi(g)^{4}}\sum_{k=1}^{d}\pi(|\phi_{k}g|^{2})\,.$
We now compute explicitly each of the quantities in the right-hand-side above using Proposition 11. First we have
$\pi(g)$ $=(2\pi(1-e^{-2t}))^{-d/2}\int e^{-V(e^{t}(z-x))}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}\mathrm{d}x\,$,
$=e^{-td}Z_{V}p_{t}^{V}(z)\,.$
Then, the variance term reads
$m_{2}(g)$ $=(2\pi(1-e^{-2t}))^{-d/2}\int(e^{-V(e^{t}(z-x))}-e^{-td}Z_{V}p_{t}^{V}(z))^{2}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}\mathrm{d}x\,$,
$=e^{-td}Z_{2V}p_{t}^{2V}(z)-2e^{-2td}Z_{2V}Z_{V}p_{t}^{V}(z)p_{t}^{2V}(z)+(e^{-td}Z_{V})^{2}p_{t}^{V}(z)^{2}\,.$
Finally, the second order moment term reads
$\sum_{k=1}^{d}\pi(|\phi_{k}g|^{2})$ $=(2\pi(1-e^{-2t}))^{-d/2}\int\|x\|^{2}e^{-2V(e^{t}(z-x))}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}\mathrm{d}x\,$,
$=e^{-td}p_{t}^{2V}(z)\Theta(z)\,$,
where we denoted
$\Theta(z)\coloneqq Z_{2V}(1-e^{-2t})^{2}(\Delta\log(p_{t}^{2V})(z)+\|\nabla\log(p_{t}^{2V})(z)\|^{2}+d/(1-e^{-2t}))\,.$
Hence, once integrated against $p_{t}^{V}$, our lower bound reads
$\int\frac{3m_{2}(g)}{n\pi(g)^{4}}\sum_{k=1}^{d}\pi(|\phi_{k}g|^{2})p_{t}^{V}(z)\mathrm{d}z=$ $~e^{2td}\frac{Z_{2V}^{2}}{Z_{V}^{4}}\int\Theta(z)\frac{p_{t}^{2V}(z)^{2}}{p_{t}^{V}(z)^{3}}\mathrm{d}z-2e^{td}\frac{Z_{2V}^{2}}{Z_{V}^{3}}\int\Theta(z)\frac{p_{t}^{2V}(z)^{2}}{p_{t}^{V}(z)^{2}}\mathrm{d}z$
$+e^{td}\frac{Z_{2V}}{Z_{V}^{2}}\int\Theta(z)\frac{p_{t}^{2V}(z)}{p_{t}^{V}(z)}\mathrm{d}z\,.$
Combining Proposition 6 and Lemma 7, we have that the third term is always finite. For the specific case $V(x)=\frac{\|x\|^{2}}{2}$, we show that the second term also remains finite while the first diverges thus making the overall lower bound diverge. First, we recall that for this choice of potential, we have
[…]
from what we obtain
$\begin{cases}&\Theta(z)=Z_{2V}\big(\|z\|^{2}+\frac{d(1-e^{-2t})[2(1-e^{-2t})+e^{-4t}]}{2-e^{-2t}}\big)\,,\\ &\frac{p_{t}^{2V}(z)^{2}}{p_{t}^{V}(z)^{2}}\propto e^{-\frac{\|e^{-t}z\|^{2}}{2(1-e^{-2t}/2)}}\,,\\ &\frac{p_{t}^{2V}(z)^{2}}{p_{t}^{V}(z)^{3}}\propto e^{\frac{\|z\|^{2}(1-e^{-2t})}{2(1-e^{-2t}/2)}}\,.\end{cases}$
In particular, we obtain that the second term is bounded and that the first term diverges to infinity making the overall bound of Agapiou et al. [2017] vacuous.
Appendix C Proof of Proposition 4
Before starting the proof, we recall the following identities.
@@ Appendix C Proof of Proposition 4
$p_{t}^{V}(z)=\frac{e^{td}}{Z_{V}}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]\,$, (23)
that
$\nabla\log(p_{t}^{V})(z)=\frac{\mathbb{E}[-Y_{t}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}\,$, (24)
@@
$\begin{split}\nabla^{2}\log(p_{t}^{V})(z)=&\frac{\mathbb{E}[Y_{t}Y_{t}^{\top}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})^{2}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}-\frac{I_{d}}{(1-e^{-2t})}\\ &-(\nabla\log(p_{t}^{V})(z)))(\nabla\log(p_{t}^{V})(z))^{\top}\,.\end{split}$ (25)
@@
$\nabla\log(p_{t}^{V})(z)=\frac{\int_{\mathbb{R}^{d}}-(z-xe^{-t})e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}{(1-e^{-2t})\int_{\mathbb{R}^{d}}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}\,.$
@@
$\nabla^{2}\log(p_{t}^{V})(z)$ $=-\frac{I_{d}}{(1-e^{-2t})}+\frac{\int_{\mathbb{R}^{d}}(z-xe^{-t})(z-xe^{-t})^{\top}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}{(1-e^{-2t})^{2}\int_{\mathbb{R}^{d}}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}$
$~~~-\frac{(\int_{\mathbb{R}^{d}}(z-xe^{-t})e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx)(\int_{\mathbb{R}^{d}}(z-xe^{-t})e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx)^{\top}}{(1-e^{-2t})^{2}(\int_{\mathbb{R}^{d}}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx)^{2}}$
$=-\frac{I_{d}}{(1-e^{-2t})}+\frac{\int_{\mathbb{R}^{d}}(z-xe^{-t})(z-xe^{-t})^{\top}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}{(1-e^{-2t})^{2}\int_{\mathbb{R}^{d}}e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}dx}$
@@
For the rest of the proof we shall drop the dependence in $z$ and write $\hat{s}_{t,n}(z)=\frac{\hat{N}}{\hat{D}}$ with the empirical numerator $\hat{N}=-\frac{\sum_{i=1}^{n}y_{i}e^{-V(e^{t}(z-y_{i}))}}{1-e^{-2t}}$ and denominator $\hat{D}=\sum_{i=1}^{n}e^{-V(e^{t}(z-y_{i}))}$ where we recall $y_{i}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$. In the following proposition, we explicitly compute the variances of $\hat{N}$ and $\hat{D}$ using the formulas above. In what follows, we shall denote $N=\mathbb{E}[N]$ and $D=\mathbb{E}[D]$
Proposition 12 (Variance of estimators)
Let $y_{1},\dots,y_{n}$ i.i.d. distributed as $\mathcal{N}(0,(1-e^{-2t})I_{d})$. Denote by $\pi$ a standard normal density, and by $\hat{N}(z)$ and $\hat{D}(z)$ the numerator and denominator of the estimator defined in Eq. 5. We have:
$\mathbb{E}[\|\hat{N}-N\|^{2}]$ $\leq\frac{p_{t}^{2V}Z_{2V}e^{-td}}{n}\left(\Delta\log(p_{t}^{2V})(z)-\frac{\Delta\log(\pi)(z)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}(z)\right)\,$,
$\mathbb{E}[\|\hat{D}-D\|^{2}]$ $\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$
@@
$\mathbb{E}[\|\hat{N}-N\|^{2}]=\frac{1}{n}\mathbb{E}\left[\left\|\frac{Y_{t}e^{-V((z-Y_{t})e^{t})}}{1-e^{-2t}}-N(z)\right\|^{2}\right]\leq\frac{1}{n}\mathbb{E}\left[\frac{\|Y_{t}\|^{2}e^{-2V((z-Y_{t})e^{t})}}{(1-e^{-2t})^{2}}\right]\,.$
@@
$\mathbb{E}[(\hat{D}-D)^{2}]=\frac{1}{n}\left(\mathbb{E}[e^{-2V((z-Y_{t})e^{t})}]-\mathbb{E}[e^{-2((z-Y_{t})e^{t})}]^{2}\right)\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$ (26)
We can now prove Proposition 4.
@@
Define the event $A=(\hat{D}\geq\eta D)\cap(\|\hat{N}\|\leq\kappa\|N\|)$ where $\eta\leq 1,\kappa\geq 1$ are positive scalars to be chosen later. We start to decompose the quadratic error as:
@@
[…]
@@
[…]
$~~~+\left(3\frac{\|\hat{N}_{\hat{t}}\|^{2}}{\hat{D}_{\hat{t}}^{2}}-2\langle\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}},\frac{N}{D}\rangle\right)(\hat{D}-D)^{2}\bigg)$
@@
Hence, almost surely over AA
@@
[…]
@@
[…]
@@
We now upper bound the probability of the event $\bar{A}=(\hat{D}<\eta D)\cup(\|\hat{N}\|>\kappa\|N\|)$. By Chebyshev’s inequality, using $\eta<1$, it holds that
@@
[…]
@@
Combining this with the bound Eq. C eventually yields